What is the speed of a plane the travels south from Edmonton to Calgary (310 km) in 30.0
min? Answer in km/hr. Place your answer in the box to the correct number of significant
digits.
1
Expert's answer
2020-05-18T10:24:50-0400
"v=\frac{s}{t}=\frac{310}{0.5}=620\frac{km}{hr}"